0.5M KOH is one of the reagents needed in the qualitative determination of stimulant cathartic anthraquinone in plants. How many grams of KOH are required to prepare 1L solution? MW KOH = 56 g/mol
1
Expert's answer
2020-10-08T13:57:14-0400

"C=\frac{n}{V}"

"n=C\times V=0.5\times1=0.5" mol


"m=n\times M=0.5\times56=28"  g